# Energy-Efficient Flying LoRa Gateways: A Multi-Agent Reinforcement Learning Approach

### v2

@@ Abstract
As next-generation Internet of Things (NG-IoT) networks continue to grow, the number of connected devices is rapidly increasing, along with their energy demands. This creates challenges for resource management and sustainability. Energy-efficient communication, particularly for power-limited IoT devices, is therefore a key research focus. In this paper, we deployed flying LoRa gateways mounted on unmanned aerial vehicles (UAVs) to collect data from LoRa end devices and transmit it to a central server. Our primary objective is to maximize the global system energy efficiency of wireless LoRa networks by joint optimization of transmission power, spreading factor, bandwidth, and user association. To solve this challenging problem, we model the problem as a partially observable Markov decision process (POMDP), where each flying LoRa GW acts as a learning agent using a cooperative multi-agent reinforcement learning (MARL). Simulation results demonstrate that our proposed method, based on the multi-agent proximal policy optimization (MAPPO) algorithm, significantly improves the global system EE and surpasses the conventional MARL schemes.
@@
Internet of Things (IoT), Long rang (LoRa), Energy efficiency, UAV communication, resource allocation.
@@ I Introduction
The next-generation Internet of Things (NG-IoT) technologies for 5G and 6G applications are revolutionizing communication by enabling seamless data exchange between devices and networks. This evolution drives intelligent applications in areas such as healthcare, smart cities, agriculture, and autonomous vehicles [1]. With IoT device connections expected to reach 125 billion by 2030 [2], energy consumption presents a significant challenge, particularly in maintaining low-power and long-range communication networks. Enhancing energy efficiency (EE) is therefore essential to align with global sustainability objectives, including the United Nations Sustainable Development Goal 7: Affordable and Clean Energy [3].
@@
Beyond infrastructure deployment, effective resource allocation plays a crucial role in optimizing LoRa network performance. Existing studies primarily rely on alternative optimization techniques, where complex optimization problems are decomposed into sub-problems and solved iteratively [4]. Although alternative optimization methods have proven effective in certain network environments, they often struggle to adapt dynamically to varying IoT traffic patterns and environmental conditions, leading to suboptimal resource utilization. For example, the authors in [5] employ an alternative optimization-based method for a single-flying LoRa GW, which lacks adaptability in dynamic environments. In contrast, reinforcement learning (RL)-based approaches, such as those in [6, 7], utilize the Q-learning approach for resource allocation. However, these methods depend on static Q-tables, making them impractical for managing complex and dynamic IoT environments. To improve the EE system of LoRa, the work in [8] proposes a framework based on deep reinforcement learning (DRL) proximal policy optimization (PPO), but it is limited to a single GW for data collection, which restricts scalability and flexibility.
To address these limitations, we propose a UAV-assisted multi-flying LoRa GW deployment. This novel approach can dynamically reposition GWs to enhance network coverage, mitigate NLoS issues, optimize end-devices (EDs) association, minimize transmission power, and maximize energy-efficient communication. Furthermore, we utilize a multi-agent reinforcement learning (MARL) framework with multi-agent PPO (MAPPO) to optimize resource allocation in multi-flying LoRa GWs. Although single-agent RL approaches have shown promise in simpler scenarios [8], extending them to multi-UAV LoRa networks introduces significant challenges, including partial observability, decentralized control, and the need to optimize multiple interdependent parameters in dynamic environments. Our MAPPO approach overcomes these challenges by allowing UAV-mounted GWs to autonomously manage key network parameters, including spreading factor (SF) allocation, transmission power (TP) control, bandwidth (W) selection, and ED association. By incorporating air-to-ground (A2G) link characteristics, our framework ensures efficient resource distribution while maximizing system-wide EE. Unlike traditional methods, MAPPO continuously updates policies in real time, making it more adaptable to dynamic network conditions. To the best of the authors’ knowledge, this is the first work to propose a MARL-based approach for optimizing global system EE in a multi-flying LoRa GW deployment. The main contributions of this paper are summarized as follows:
  • We formulate the resource allocation and user association problem as a single optimization problem aimed at maximizing EE in LoRa networks, considering both device positions and A2G propagation.
@@
  • We leverage the MARL-based MAPPO approach that enables flying LoRa GWs to learn a centralized value function while optimizing their policies to maximize global system EE.
  • Our proposed framework leverages one-to-many matching algorithms for an efficient device-to-gateway association, demonstrating faster and more stable convergence compared to state-of-the-art MARL algorithms.
@@ II System Model
We consider a resource allocation framework for uplink transmission in a LoRa network consisting of $\mathcal{V}$ LoRa EDs, $\mathcal{U}$ flying GWs, and a single network server, as shown in Fig. 1. In this setup, UAVs equipped with LoRa GWs are deployed over a square target area $S$. Each flying LoRa GW has a limited communication range and can simultaneously connect to multiple EDs within its association quota $\Lambda_{max}$. Specifically, each GW collects and decodes packets from all eligible EDs in range and relays these packets to the network server. The sets of EDs and GWs are denoted by $\mathcal{V}=\{1,\ldots,V\}$ and $\mathcal{U}=\{1,\ldots,U\}$, respectively.
@@ II-B Air-to-Ground Channel Model
In UAV communication systems, A2G link propagation plays a crucial role and can be leveraged for flying GWs communication [9]. In this work, a probabilistic path loss model was used to model A2G communication links, whereby the line-of-sight (LoS) and NLoS links were considered separately with different probabilities of occurrence. Consequently, the likelihood of having a LoS connection between GW $u$, and an ED $v$ is given by:
$P_{LoS}(\hat{\phi})=\frac{1}{1+\vartheta e^{-\lambda(\hat{\phi}-\vartheta)}},$ (1)
where $\hat{\phi}$ = $sin^{-1}(\frac{h_{u}}{d_{u,v}})$ is the angle of elevation from GWs $u$, $u\in\mathcal{U}$ to EDs $v$, $v\in\mathcal{V}$. $h_{u}$ is the altitude of the UAV $u$. The Euclidean distance between $u$ and $v$ is denoted as $d_{u,v}$. $\vartheta$ and $\lambda$ are constants that depend on the environment. Note, from Eq. (1), the probability values for $P_{NLoS}$ can be expressed as $P_{NLoS}(\hat{\phi})=1-P_{LoS}(\hat{\phi})$. Furthermore, the path loss model from UAV $u$ to the ground LoRa ED $v$ can be expressed as:
$L^{LoS}_{u,v}$ $=L^{FS}(d_{r})+10\delta_{LoS}d_{u,v}+\mathcal{X}\sigma_{LoS},$ (2a)
$L^{NLoS}_{u,v}$ $=L^{FS}(d_{r})+10\delta_{NLoS}d_{u,v}+\mathcal{X}\sigma_{NLoS},$ (2b)
where $L^{FS}$ denotes the free-space path-loss with $L^{FS}(d_{r})=20\log_{10}\left(\frac{4\pi d_{r}f}{c}\right)$. $\delta$ is the path loss exponent. $f$ is the carrier frequency in Hz, $c$ in m/s is the speed of light, and $d_{r}$ is the reference distance. $\mathcal{X}\sigma_{LoS}$ and $\mathcal{X}\sigma_{NLoS}$ are the shadowing random variables which are characterized as Gaussian random variables with zero mean and $\sigma_{LoS}$ and $\sigma_{NLoS}$ standard deviations. Consequently, the overall A2G path loss between GW $u$ and LoRa ED $v$ is characterized as:
$l_{u,v}^{a2g}=P_{LoS}(\hat{\phi})L^{LoS}_{u,v}+P_{NLoS}(\hat{\phi})L^{NLoS}_{u,v}.$ (3)
@@ II-C Energy Efficiency Model
To model the EE of our proposed system, we begin by linking the network topology to the LoRa PHY parameters. This is achieved by first modeling the signal-to-noise ratio (SNR) between UAV $u$ and ED $v$ at time slot $t$, expressed as
$\rho_{u,v}(t)=\frac{P_{v}(t)\cdot G_{u,v}}{\sigma^{2}},$ (4)
where $G_{u,v}=10^{-{l_{u,v}^{\text{a2g}}}/{10}}$ represents the channel gain, $P_{v}(t)$ is the TP of an ED at time $t$, and $\sigma^{2}$ is the noise power. Building on this, we calculate the signal-to-interference-plus-noise ratio (SINR) $\mho_{u,v}^{n}(t)$ between UAV $u$ and ED $v$ using the $n$-th SF at time slot $t$ as follows:
@@
where $W_{v}(t)$ is the bandwidth allocated to the communication link. Furthermore, the EE of each UAV is calculated by dividing the sum of all uplink data rates from all EDs connected to the UAV by the total consumed power. Hence, we define the EE $\zeta_{u}(t)$ of UAV $u$ at time $t$ as:
@@ III Problem Formulation
This section provides a detailed description of the optimization problem based on our proposed system model. The objective of this paper is to maximize the global system EE of flying LoRa GWs, as defined in Eq. (11a). Consequently, we formulate our optimization problem as follows:
@@ IV Multi-Agent PPO Algorithm
2) States $\mathcal{S}$: The global state represents the complete environment configuration at time step $t$. Therefore, we define the global state vector of all UAVs as $\mathbf{\mathcal{S}}(t)=\left\{s_{1}(t),s_{2}(t),\dots,s_{u}(t),\dots,s_{U}(t)\right\}$, where $s_{u}(t)$ denotes the local state of agent $u\in\mathcal{U}$ at time $t$.
3) Observations $\mathcal{O}$: Due to partial observability, each flying LoRa GW $u$ at time $t$ perceives only a subset of the global state $\mathbf{\mathcal{S}}(t)$. Hence, the local observation $o_{u}(t)\in\mathcal{O}_{u}\subset\mathcal{S}(t)$ for agent $u$ is denoted as $o_{u}(t)=\left\{\psi(t),p(t),\Re(t),assEDs_{pos}(t),GWs_{pos}(t)\right\}$. Here, $assEDs_{pos}(t)$ represents the positions of the EDs currently associated with GW $u$, and $GWs_{pos}(t)$ are the positions of neighboring GWs within communication range. This partial observation mechanism ensures that each GW makes decisions based on its local view of the environment.
TABLE I: SINR Thresholds with $W=125$ kHz [6]
SF 7 8 9 10 11 12
SNRthreshold (dB) -7.5 -10 -12.5 -15 -17.5 -20
4) Actions $\mathcal{A}$: Each agent $u\in\mathcal{U}$ selects an action $a_{u}(t)\in\mathcal{A}_{u}$ at time $t$. The action comprises the selection of communication parameters including spreading factor, transmission power, and bandwidth. This can be represented as $a_{u}(t)=\left\{\psi_{u}(t),p_{u}(t),w_{u}(t)\right\}$, where $\psi_{u}(t)$, $p_{u}(t)$, and $w_{u}(t)$ denote the SF, TP, and bandwidth chosen by agent $u$ at time $t$, respectively.
5) Reward Function $\mathcal{R}$: The reward for each flying LoRa GW $u$ at time step $t$ is defined as:
@@
Note that the reward in Eq.(12) is allocated only if the SNR constraint (11h) is satisfied; otherwise, the reward is zero. This constraint can be formulated as:
[…]
Therefore, the cumulative reward over the entire time horizon $T$ and all GWs $\mathcal{U}$ is:
$\mathcal{R}=\sum_{t=1}^{T}\sum_{u=1}^{U}r_{u}(t)$ (14)
6) Policy function: The policy function $\boldsymbol{\pi_{\theta_{u}}}(\textbf{a}_{u},o_{u})$ is modeled by an actor-network, with the vector $\boldsymbol{\theta}_{u}$ as its parameters. It determines the strategy for the flying GWs based on their local observations.
7) Value function: The value function $V_{\phi}(\boldsymbol{o_{u}})$ is implemented by a critic network parameterized by $\phi$. This network evaluates the expected future rewards for flying GWs given the current state $\boldsymbol{o}_{u}$. The critic aims to learn a value function that approximates the optimal future rewards, guiding the agents toward the globally optimal policy that maximizes their long-term rewards.
@@ V-A Simulation setup
To satisfy the LoRa quality of service constraint, probabilistic path loss exponents for LoS and NLoS are given by $\delta_{\text{LoS}}=2$ and $\delta_{\text{NLoS}}=2.5$. We consider a Gaussian random variables $\mathcal{X}\sigma_{\text{LoS}}=5$ and $\mathcal{X}\sigma_{\text{NLoS}}=20$. In addtion, we consider a carrier frequency of $f=868\,\text{MHz}$. Furthermore, the agents can dynamically select configurations from three adjustable parameters. Particularly, SF is chosen from $\{7,8,9,10,11,12\}$, TP is selected from $\{2,5,8,11,14\}\,\text{dBm}$, and bandwidth $W$ configured from $\{125,250,500\}\,\text{kHz}$. For training, we employ a recurrent neural network (RNN) with 128 hidden units, a learning rate of $\alpha=3\times 10^{-4}$ for both actor and critic networks, and a soft target update coefficient of $0.01$. The training protocol uses a PPO clip range of $\epsilon=0.2$, a discount factor of $\gamma=0.99$, and the Adam optimizer.
In our simulation, we compare our proposed approach with three other MARL algorithms: Counterfactual Multi-Agent (COMA) [14], Multi-Agent Advantage Actor-Critic (MAA2C) [15], and Value-Decomposition Networks (VDN)[16]. The COMA algorithm is an actor-critic method in which a centralized critic estimates the Q-function, with decentralized actors optimizing their policies. The MAA2C algorithm extends the A2C algorithm to multi-agent scenarios by incorporating a centralized critic that learns the joint value function. At the same time, each agent maintains its own actor to learn individual policies. Finally, the VDN framework is designed for cooperative MARL tasks, wherein each agent learns a distinct value function that is decomposed into shared and local value functions.

### v3

@@ Abstract
As next-generation Internet of Things (NG-IoT) networks continue to grow, the number of connected devices is rapidly increasing, along with their energy demands. This creates challenges for resource management and sustainability. Energy-efficient communication, particularly for power-limited IoT devices, is therefore a key research focus. In this paper, we deployed flying LoRa gateways mounted on unmanned aerial vehicles (UAVs) to collect data from LoRa end devices and transmit it to a central server. Our primary objective is to maximize the global system energy efficiency of wireless LoRa networks by joint optimization of transmission power, spreading factor, bandwidth, and user association. To solve this challenging problem, we model the problem as a partially observable Markov decision process (POMDP), where each flying LoRa GW acts as a learning agent using a cooperative multi-agent reinforcement learning (MARL). Simulation results demonstrate that our proposed method, based on the multi-agent proximal policy optimization algorithm, significantly improves the global system energy efficiency and surpasses the popular MARL and other conventional schemes.
@@
Internet of Things (IoT), Long Rang (LoRa), Energy efficiency, UAV, Resource allocation.
@@ I Introduction
The next-generation Internet of Things (NG-IoT) technologies for 5G and 6G applications are revolutionizing communication by enabling seamless data exchange between devices and networks. This evolution drives intelligent applications in areas such as healthcare, smart cities, agriculture, and autonomous vehicles [1]. With IoT device connections expected to reach 125 billion by 2030 [2], energy consumption presents a significant challenge, particularly in maintaining low-power and long-range communication networks. Enhancing energy efficiency is therefore essential to align with global sustainability objectives, including the United Nations Sustainable Development Goal 7: Affordable and Clean Energy [3].
@@
Beyond infrastructure deployment, effective resource allocation plays a crucial role in optimizing LoRa network performance. Existing studies primarily rely on alternative optimization techniques, where complex optimization problems are decomposed into sub-problems and solved iteratively [4]. Although alternative optimization methods have proven effective in certain network environments, they often struggle to adapt dynamically to varying IoT traffic patterns and environmental conditions, leading to suboptimal resource utilization. For example, the authors in [5] employ an alternative optimization-based method for a single-flying LoRa GW, which lacks adaptability in dynamic environments. In contrast, reinforcement learning (RL)-based approaches, such as those in [6, 7], utilize the Q-learning approach for resource allocation. However, these methods depend on static Q-tables, making them impractical for managing complex and dynamic IoT environments. To improve the energy efficiency system of LoRa, the work in [8] proposes a framework based on deep reinforcement learning (DRL) proximal policy optimization (PPO), but it is limited to a single GW for data collection, which restricts scalability and flexibility.
To address these limitations, we propose a UAV-assisted multi-flying LoRa GW deployment. This novel approach can dynamically reposition GWs to enhance network coverage, mitigate NLoS issues, optimize end-devices (EDs) association, minimize transmission power, and maximize energy-efficient communication. Furthermore, we utilize a multi-agent reinforcement learning (MARL) framework with multi-agent PPO (MAPPO) to optimize resource allocation in multi-flying LoRa GWs. Although single-agent RL approaches have shown promise in simpler scenarios [8], extending them to multi-UAV LoRa networks introduces significant challenges, including partial observability, decentralized control, and the need to optimize multiple interdependent parameters in dynamic environments. Our MAPPO approach overcomes these challenges by allowing UAV-mounted GWs to autonomously manage key network parameters, including spreading factor (SF) allocation, transmission power (TP) control, bandwidth (W) selection, and ED association. By incorporating air-to-ground (A2G) link characteristics, our framework ensures efficient resource distribution while maximizing system-wide energy efficiency. Unlike traditional methods, MAPPO continuously updates policies in real time, making it more adaptable to dynamic network conditions. To the best of the authors’ knowledge, this is the first work to propose a MARL-based approach for optimizing global system energy efficiency in a multi-flying LoRa GW deployment. The main contributions of this paper are summarized as follows:
  • We formulate the resource allocation and user association problem as a single optimization problem aimed at maximizing energy efficiency in LoRa networks, considering both device positions and A2G propagation.
@@
  • We leverage the MARL-based MAPPO approach that enables flying LoRa GWs to learn a centralized value function while optimizing their policies to maximize global system energy efficiency.
  • Our proposed framework leverages one-to-many matching algorithms for an efficient device-to-gateway association, demonstrating faster and more stable convergence compared to state-of-the-art MARL algorithms and other conventional allocation schemes.
@@ II System Model
We consider a resource allocation framework for uplink transmission in a LoRa network consisting of $\mathcal{V}$ LoRa EDs, $\mathcal{U}$ flying GWs, and a single network server, as shown in Fig. 1. In this setup, UAVs equipped with LoRa GWs are deployed over a square target area $S$. Each flying LoRa GW has a limited communication range and can simultaneously connect to multiple EDs within its association quota $\Lambda_{max}$. Specifically, each GW collects and decodes packets from all associated EDs in range and relays these packets to the network server. The sets of EDs and GWs are denoted by $\mathcal{V}=\{1,\ldots,V\}$ and $\mathcal{U}=\{1,\ldots,U\}$, respectively.
@@ II-B Air-to-Ground Channel Model
In this study, we adopt a probabilistic path-loss model for A2G links, accounting separately for Line-of-Sight (LoS) and NLoS propagation scenarios based on their occurrence probabilities. Consequently, the probability of having a LoS connection between GW $u\in\mathcal{U}$ and an ED $v\in\mathcal{V}$ is computed as:
$P_{\text{LoS}}(\hat{\phi})=\frac{1}{1+\alpha e^{-\lambda(\hat{\phi}-\alpha)}},$ (1)
where $\hat{\phi}=\tan^{-1}\left(\frac{h_{u}}{d_{u,v}}\right)$ represents the elevation angle between UAV $u$ at altitude $h_{u}$ and ED $v$. The horizontal Euclidean distance $d_{u,v}$ between UAV GW $u$ and ED $v$ is calculated based on their Cartesian coordinates as $d_{u,v}=\sqrt{(x_{u}-x_{v})^{2}+(y_{u}-y_{v})^{2}},$ where $(x_{u},y_{u})$ and $(x_{v},y_{v})$ represent the horizontal positions of UAV $u$ and ED $v$, respectively. Here, the constants $\alpha$ and $\lambda$ are determined empirically based on environmental conditions. Note, from Eq.(1), the probability of the NLoS can be expressed as $P_{\text{NLoS}}(\hat{\phi})=1-P_{\text{LoS}}(\hat{\phi})$.
Furthermore, we consider a free-space path loss (FSPL) at distance $d_{u,v}$ with carrier frequency $f$ as $L^{\text{FS}}(d_{u,v})=20\log_{10}\left(\frac{4\pi d_{u,v}f}{c}\right).$ Here, $c$ is the speed of light. Finally, we model the overall A2G path loss according to [9], by incorporating both LoS and NLoS conditions given as:
$L_{u,v}^{\text{A2G}}=L^{\text{FS}}(d_{u,v})+\eta_{\text{LoS}}P_{\text{LoS}}(\hat{\phi})+\eta_{\text{NLoS}}P_{\text{NLoS}}(\hat{\phi}),$ (2)
where $\eta_{\text{LoS}}$ and $\eta_{\text{NLoS}}$ are additional loss factors associated with LoS and NLoS conditions, respectively.
@@ II-C Energy Efficiency Model
To model the energy efficiency of our proposed system, we begin by linking the network topology to the LoRa PHY parameters. This is achieved by first modeling the signal-to-noise ratio (SNR) between UAV $u$ and ED $v$ at time slot $t$, expressed as $\rho_{u,v}(t)=\frac{P_{v}(t)\cdot G_{u,v}}{\sigma^{2}}$, where $G_{u,v}=10^{-{L_{u,v}^{\text{A2G}}}/{10}}$ represents the channel gain, $P_{v}(t)$ is the TP of an ED at time $t$, and $\sigma^{2}$ is the noise power. Building on this, we calculate the signal-to-interference-plus-noise ratio (SINR) $\mho_{u,v}^{n}(t)$ between UAV $u$ and ED $v$ using the $n$-th SF at time slot $t$ as follows:
@@
where $W_{v}(t)$ is the bandwidth allocated to the communication link. Furthermore, the energy efficiency of each UAV is calculated by dividing the sum of all uplink data rates from all EDs connected to the UAV by the total consumed power. Hence, we define the energy efficiency $\zeta_{u}(t)$ of UAV $u$ at time $t$ as:
@@ III Problem Formulation
This section provides a detailed description of the optimization problem based on our proposed system model. The objective of this paper is to maximize the global system energy efficiency of flying LoRa GWs, as defined in Eq. (9a). Consequently, we formulate our optimization problem as follows:
@@ IV Multi-Agent PPO Algorithm
2) Observations $\mathcal{O}$: Due to partial observability, each flying LoRa GW $u$ at time $t$ perceives only a subset of the global state $\mathbf{\mathcal{S}}(t)$. Hence, the local observation $o_{u}(t)\in\mathcal{O}_{u}\subset\mathcal{S}(t)$ for agent $u$ is denoted as $o_{u}(t)=\left\{\psi(t),p(t),\Re(t),assEDs_{pos}(t),GWs_{pos}(t)\right\}$. Here, $assEDs_{pos}(t)$ represents the positions of the EDs currently associated with GW $u$, and $GWs_{pos}(t)$ are the positions of neighboring GWs within communication range. This partial observation mechanism ensures that each GW makes decisions based on its local view of the environment.
3) Actions $\mathcal{A}$: Each agent $u\in\mathcal{U}$ selects an action $a_{u}(t)\in\mathcal{A}_{u}$ at time $t$. The action comprises the selection of communication parameters including spreading factor, transmission power, and bandwidth. This can be represented as $a_{u}(t)=\left\{\psi_{u}(t),p_{u}(t),w_{u}(t)\right\}$, where $\psi_{u}(t)$, $p_{u}(t)$, and $w_{u}(t)$ denote the SF, TP, and W chosen by agent $u$ at time $t$, respectively.
4) Reward Function $\mathcal{R}$: The reward for each flying LoRa GW $u$ at time step $t$ is defined as:
@@
Note that the reward in Eq.(10) is allocated only if the SNR constraint (9h) is satisfied; otherwise, the reward is zero. Therefore, the cumulative reward over the entire time horizon $T$ and all GWs $\mathcal{U}$ is:
$\mathcal{R}=\sum_{t=1}^{T}\sum_{u=1}^{U}\gamma^{t}r_{u}(t).$ (11)
@@ V-A Simulation setup
In our simulations, the LoS and NLoS path loss coefficients are set as $\eta_{\text{LoS}}=0.1$ and $\eta_{\text{NLoS}}=21$, respectively. Additionally, we assume the path loss exponent factor $\lambda=0.43$, and the environment-dependent constant $\alpha=4.88$ [9]. Furthermore, we consider a carrier frequency of $f=868\,\text{MHz}$, the speed of light $c=3\times 10^{8}m/s$, and the noise power is set to $\sigma^{2}=10^{-15}W$. Furthermore, the agents can dynamically select configurations from three adjustable parameters. Particularly, SF is chosen from $\{7,8,9,10,11,12\}$, TP is selected from $\{2,5,8,11,14\}\,\text{dBm}$, and bandwidth $W$ configured from $\{125,250,500\}\,\text{kHz}$.
To train our agents, we employ a neural network architecture based on a gated recurrent unit (GRU) with 128 hidden units, a learning rate of $\alpha=3\times 10^{-4}$ for both actor and critic networks, and a soft target update coefficient of $0.01$. The training protocol uses a PPO clip range of $\epsilon=0.2$, a discount factor of $\gamma=0.99$, and the Adam optimizer.
As benchmarks, we evaluate our approach using two distinct schemes: the popular DRL-based algorithms and the conventional allocation schemes. For the former, we used MAA2C [14], COMA [15], and QMIX [16]. The MAA2C algorithm extends A2C to multi-agent scenarios by incorporating a centralized critic that learns the joint value function, while each agent maintains its own actor to learn individual policies. COMA is an actor-critic method where a centralized critic estimates the Q-function, and decentralized actors optimize their policies. QMIX is a deep RL method that promotes cooperative behavior by decomposing the value function in a way that preserves monotonicity. For the latter, we used a genetic-based algorithm and a distance-based allocation. Specifically, the distance-based approach assigns higher SFs and TP to more distant EDs. On the other hand, the genetic algorithm assigns different SFs, TP, and W to EDs to satisfy the SNR constraint in (9h). This algorithm is configured with a population size of 50, an elite size of 5, a mutation rate of 0.1, and 300 generations.